OpenStudy (anonymous):

Calculate the unit vector (in terms of i(hat) and j(hat) ) in the direction opposite to the direction of each of the following vectors. Vector A= 3.7i(hat) + 4.5j(hat)

OpenStudy (ash2326):

@Bakken2012 could you find the unit vector for A?

OpenStudy (anonymous):

no im confused on how to do it

OpenStudy (ash2326):

Ok, I'll explain you

OpenStudy (ash2326):

Unit vector is a vector with magnitude 1 suppose A' is a unit vector \[A'=ai+bj\] then \[\sqrt{a^2+b^2}=1\] If we have a vector B=ci+dj B is not a unit vector, but a unit vector in direction of b would be \[\frac{ci}{\sqrt{c^2+d^2}} \ \ \ \ \ \ \ \ \ \ +\frac{dj}{\sqrt{c^2+d^2}}\] do you understand this?

OpenStudy (anonymous):

ok i understand that, but i beleive I am supposed to calculate it in the opposite direction

OpenStudy (ash2326):

Yeah, unit vector in opposite direction is \[-\frac{ci}{\sqrt{c^2+d^2}} \ \ \ \ \ \ \ \ \ \ -\frac{dj}{\sqrt{c^2+d^2}}\]

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

@Bakken2012 could you do it now?

OpenStudy (anonymous):

i think so let me try

OpenStudy (ash2326):

cool :)

OpenStudy (phi):

in pictures: